OpenStudy (anonymous):

What is the value of x in the solution to the system of linear equations? y=3x+2 y=x-4

OpenStudy (anonymous):

simply equate both eq as both are equal y so you get x-4=3x+2 =>3x-x=-4-2 =>2x=-6 =>x=-3

OpenStudy (anonymous):

so since you have 2 equations both equal to y we know they can be equal to each other soooo..... 3x+2=x-4 2x+2=-4 2x=-6 x=-3 then plug that in into any of the equations to get y. I chose y=x-4 just cuz . y=(-3)-4 y=-7 so your final solution will be x=-3 and y=-7 (-3,-7)